OpenStudy (anonymous):

Can someone assist me with this curvature problem: Compute the curvature and the principal unit normal of the ellpitical hellix described by the equations: x=cost, y=2sint, and z=t

OpenStudy (amistre64):

is curvature gonna be binormal as well?

OpenStudy (amistre64):

or just a k

OpenStudy (anonymous):

we are trying to find k or curvature

OpenStudy (amistre64):

\[k=\frac{|r'xr''|}{|r'|^3}\] if i recall it correctly

OpenStudy (anonymous):

or 1/magnitude of V *magniutde of Tangent unit vector

OpenStudy (anonymous):

but your formula looks cool, so lets try taht one please

OpenStudy (amistre64):

r = <x,y,z> =<cos(t), 2sin(t), t> r' = <-sin(t), 2cos(t),1> r'' = <-cos(t), -2sin(t),0>

OpenStudy (anonymous):

okay, okay, that what i got...

OpenStudy (amistre64):

r' x r'' x -sin(t) -cos(t) x=2sin(t) = 2sin(t) y 2cos(t) -2sin(t) -y=cos(t) = -cos(t) z 1 0 z=2sin^2(t) + 2cos^2(t) = 2 r' x r'' = <2sin(t), -cos(t),2> whats our magnitude? of that?

OpenStudy (anonymous):

so did you cross r prime and r double prime

OpenStudy (amistre64):

i hope so :)

OpenStudy (anonymous):

hmm, , so the magnitude of prime is..

OpenStudy (anonymous):

sqrt(sin^2t+4cos^2t)

OpenStudy (anonymous):

now how do we simplify that

OpenStudy (amistre64):

|r'| = |<-sin(t), 2cos(t),1>| = sqrt(sin^2(t)+4cos^2(t)+1)

OpenStudy (anonymous):

oh i missed that 1, but anyway how do we simplify that

OpenStudy (amistre64):

well, sin^2 = 1-cos^2 sqrt(sin^2(t)+4cos^2(t)+1) sqrt(1-cos^2(t)+4cos^2(t)+1) sqrt(3cos^2(t)+2) is one option

OpenStudy (anonymous):

thats seems like the best option

OpenStudy (amistre64):

|r' x r''| = sqrt(4sin^2+cos^2+4) sooo, if my memory serves me .... IF ..... \[\frac{\sqrt{4sin^2(t)+cos^(t)+4}}{(\sqrt{3cos^2(t)+2})^3}\] is something to play with

OpenStudy (anonymous):

oh we take the magnitude of the cross product as well

OpenStudy (anonymous):

thats seems impossible to simplify

OpenStudy (amistre64):

\[\frac{\sqrt{4-4cos^2(t)+cos^2(t)+4}}{(\sqrt{3cos^2(t)+2})^3}\] \[\frac{\sqrt{-3cos^2(t)+8}}{(3cos^2(t)+2) \sqrt{3cos^2(t)+2}}\]

OpenStudy (anonymous):

what happened to the denominator how did you factor that?

OpenStudy (amistre64):

\[(\sqrt{a})^3=\sqrt{a}\sqrt{a}\sqrt{a}=a\sqrt{a}\]

OpenStudy (anonymous):

ah, yes good man

OpenStudy (amistre64):

now the radical covers top to bottom and we might be able to play inside it

OpenStudy (anonymous):

do you have a fancy formula for computing the principal unit normal?

OpenStudy (amistre64):

r'' is the normal; so just unit it

OpenStudy (amistre64):

r''/|r''|

OpenStudy (anonymous):

reall????thats is krazy

OpenStudy (anonymous):

are you doubly sure?

OpenStudy (amistre64):

yep

OpenStudy (amistre64):

if you dot r' and r''; are they perped?

OpenStudy (anonymous):

they would have to =0

OpenStudy (amistre64):

r' = <-sin(t), 2cos(t),1> r'' = <-cos(t), -2sin(t),0> ---------------------- sincos -4sincos +0 great, now I gotta wonder lol hold on

OpenStudy (anonymous):

yeah, man, this is going haywire

OpenStudy (amistre64):

http://tutorial.math.lamar.edu/Classes/CalcIII/Curvature.aspx so far so good

OpenStudy (amistre64):

lets finish up k, then well worry about unit normals ...

OpenStudy (anonymous):

okay, yeah lets finish up K, at least i need that cause then i can try and find the pricipal

OpenStudy (amistre64):

http://www.wolframalpha.com/input/?i=sqrt%284sin%5E2%28t%29%2Bcos%28t%29%2B4%29%2F%28sqrt%283cos%5E2%28t%29%2B2%29%5E3+ unless we messed up a math someplse this looks to be it

OpenStudy (anonymous):

so, we are saying thats the curvature

OpenStudy (amistre64):

yes, but im checking the math with the wolf first :)

OpenStudy (amistre64):

as far as I can tell, thats good

OpenStudy (anonymous):

alright,so then for the principal unit......

OpenStudy (amistre64):

r'/|r'| = unit T T'/|T'| = unit N

OpenStudy (amistre64):

since T' = r'' /////

OpenStudy (amistre64):

\[N=\frac{r''}{|r''|}=\frac{1}{\sqrt{cos^2+4sin^2}}<-cos(t), -2sin(t),0> \] \[N=\frac{r''}{|r''|}=\frac{1}{\sqrt{4-3cos^2}}<-cos(t), -2sin(t),0> \]

OpenStudy (anonymous):

thanks for your help peppy

OpenStudy (amistre64):

i dont spose thats a common little nickname all the young kids are throwing about these days :)

OpenStudy (anonymous):

no but seriously thanks for helping me, my teacher will be so happy,now i have to go write my bookreport on Dr. Seuss

OpenStudy (amistre64):

I hear the Lorax is in theatres ;)

OpenStudy (amistre64):

make sure to go thru the math and adjust for errors that pervade

OpenStudy (anonymous):

sure thing

OpenStudy (anonymous):

you still here?

OpenStudy (anonymous):

witch one is the principla unit normal, i think you wrote down two of them?

OpenStudy (anonymous):

@amistre64

OpenStudy (amistre64):

id have to read up on what defines a principle; im sure its the one that heads into the curve ....

OpenStudy (amistre64):

http://home.iitk.ac.in/~psraj/mth101/lecture_notes/lecture25.pdf the principle Normal is defined as T'\|T'|

OpenStudy (amistre64):

except for a usual / for division and not that \ that gets in the way :)